Compound E: C7H6O3


Hi, can someone please help me with the following questions?


Draw a molecule that combines both of the above fr

Draw a molecule that combines both of the above fr

Draw a molecule that combines both of the above fr

Draw a molecule that combines both of the above fr

Draw a molecule that combines both of the above fr

Draw a molecule that combines both of the above fr

Draw a molecule that combines both of the above fr
Answer #1

NOT in the compound E amine and nitrile
Answer #2

24)

a) Double bond equivalent can be calculated as

DBE = C - \frac{H}{2} + \frac{N}{2} +1

DBE = 7 - (6/2) + 0+1 = 5

b) due to absence of nitrogen the following groups are definitely not present

Nitrile and amine

Also, alkyne is absent

25)

The NMR signals above 5ppm shows the presence of aromatic ring

25) The singlet at 11 ppm is due to carboxylic group

25) The compound could be

он C7HsO2

25) the name of that functional group is : Alcohol (Hydroxyl group)

25) OHThe stucture of given compound E.